OpenStudy (highschoolmom2010):

answer check please? A circle is inscribed in a quadrilateral whose four angles have measures 85, 76, 94, and 105. Find the measures of the four arcs between consecutive points of tangency.

OpenStudy (highschoolmom2010):

arc 1= 180-85 arc 2=180-76 arc 3=180-94 arc 4 =180-105 arc 1= 95 arc 2=104 arc 3=86 arc 4= 75

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

good job !

OpenStudy (highschoolmom2010):

hooray thanks @ganeshie8